OpenStudy (anonymous):

A survey asks high school seniors whether they would be willing to pay $5 for their yearbook. Out if the 225 seniors surveryed,198 said "yes". What percentof the seniors said "yes"?

OpenStudy (anonymous):

how to drag the moving image http://www.mathopenref.com/trianglecircumcenter.html on powerpoint slides

OpenStudy (anonymous):

x% of 225 is 198 (x/100)*225=198 x = 198*4/9 = 88

OpenStudy (anonymous):

Thanks:)